# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 2.3 Salvetti complex
The following is a construction originated in [GP12b].
@@
Let $W_{\Gamma}$ be a finite Coxeter group with its canonical representation $\rho:W_{\Gamma}\to GL(n,\mathbb{R})$ [Bou02]. Recall a reflection in $W_{\Gamma}$ is a conjugate of a standard generator of $W_{\Gamma}$. Each reflection fixes a hyperplane in $\mathbb{R}^{n}$, which we call a reflection hyperplane. Let $\mathcal{A}$ be the collection of all reflection hyperplanes in $\mathbb{R}^{n}$. The hyperplane arrangement $\mathcal{A}$ is called the reflection arrangement associated with $W_{\Gamma}$. The following facts are standard, see e.g. [Par14].
@@
    The fundamental group $\pi_{1}(M(\mathcal{A}\otimes\mathbb{C}))$ is isomorphic to the pure Artin group $PA_{\Gamma}$ [vdL83];
@@
    As the action of $W_{\Gamma}$ permutes elements in $\mathcal{A}$, there is an induced action $W_{\Gamma}\curvearrowright M(\mathcal{A}\otimes\mathbb{C})$ and an induced action $W_{\Gamma}\curvearrowright\widehat{\Sigma}_{\mathcal{A}}$ which are free. The quotient of each of these two spaces have $\pi_{1}$ isomorphic to $A_{\Gamma}$.
@@
    The 2-skeleton of $\widehat{\Sigma}_{\mathcal{A}}/W_{\Gamma}$ is isomorphic to the presentation complex of $A_{\Gamma}$.
@@
When $\mathcal{A}$ is the reflection arrangement associated with a finite Coxeter group $W_{\Gamma}$, we will also write $\Sigma_{\mathcal{A}}$ and $\widehat{\Sigma}_{\mathcal{A}}$ as $\Sigma_{\Gamma}$ and $\widehat{\Sigma}_{\Gamma}$. Note that the 1-skeleton of $\widehat{\Sigma}_{\mathcal{A}}$ is isomorphic to the Cayley graph of $W_{\Gamma}$, and the 1-skeleton of $\Sigma_{\Gamma}$ is isomorphic to the unoriented Cayley graph of $W_{\Gamma}$ (by smashing each double edge of the usual Cayley graph to a single edge). Thus edges of $\widehat{\Sigma}_{\mathcal{A}}$ and $\Sigma_{\mathcal{A}}$ are labeled by vertices of $\Gamma$. Let $K$ be a subset, or an edge path in $\widehat{\Sigma}_{\mathcal{A}}$ or $\Sigma_{\mathcal{A}}$. Then $\operatorname{Supp}(K)$ is defined to the collection of vertices of $\Gamma$ which appear as the label of an edge which is contained in $K$. The type of a standard subcomplex of $\Sigma_{\Gamma}$ or $\widehat{\Sigma}_{\Gamma}$ is defined to be the support of this subcomplex.
@@ 2.4 Falk complexes for affine arrangements
Let $\mathcal{A}$ be an affine arrangement. Let $D_{\mathcal{A}}$ be the union of all elements in $\operatorname{Fan}(\mathcal{A})$ which are bounded in $\mathbb{C}$. Note that $D_{\mathcal{A}}$ is polyhedron complex, whose open cells are elements in $\operatorname{Fan}(\mathcal{A})$ that are contained in $D_{\mathcal{A}}$. A face of $D_{\mathcal{A}}$ is defined to be the closure of an open cell of $D_{\mathcal{A}}$. Each face of $D_{\mathcal{A}}$ is a disjoint union of fans. Top dimensional faces in $D_{\mathcal{A}}$ are in 1-1 correspondence with components of $V\setminus\cup_{H\in\mathcal{A}}H$ which are bounded. A face $F$ of $D_{\mathcal{A}}$ is dual to a face $F^{\prime}$ of $\Sigma_{\mathcal{A}}$ if the barycenter of $F^{\prime}$ (as defined in the beginning of Section 2.2) is contained in the interior of $F$. In this case, we will also say $F\subset D_{\mathcal{A}}$ is dual to the standard subcomplex $\widehat{F}^{\prime}$ of $\widehat{\Sigma}_{\mathcal{A}}$.
@@
We now define a simplex complex of group structure $\mathcal{U}$ on $D_{\mathcal{A}}$ as follows. For a face $F$ of $D_{\mathcal{A}}$, the local group at $F$ is defined to be the fundamental group of the standard subcomplex of $\widehat{\Sigma}_{\mathcal{A}}$ which is dual to $F$. The morphisms between the local groups are induced by inclusions of the associated subcomplexes. By Lemma 2.13, all the morphisms between the local groups are injective, hence $\mathcal{U}$ is a simple complex of groups. Moreover, it follows from the retraction in Definition 2.9 that each local group injects into $\pi_{1}(\widehat{\Sigma}_{\mathcal{A}})$. Thus $\mathcal{U}$ is developable with $\pi_{1}(\widehat{\Sigma}_{\mathcal{A}})=\pi_{1}\mathcal{U}$.
@@
Let $\mathbb{D}_{\mathcal{A}}$ be the development complex of $\mathcal{U}$ (cf. [BH99, Chapter II.12]). Note that $\mathbb{D}_{\mathcal{A}}$ is simply-connected. The complex $\mathbb{D}_{\mathcal{A}}$ is defined to be the Falk complex of the affine arrangement $\mathcal{A}$, motivated by the work of Falk [Fal95]. We now give an alternative description of $\mathbb{D}_{\mathcal{A}}$. Let $\widetilde{K}$ be the universal cover of $\mathbb{D}_{\mathcal{A}}$. We start with a disjoint union of a collection $\mathcal{C}$ of multiple copies of faces of $D_{\mathcal{A}}$ as follows: for each face $F$ of $D_{\mathcal{A}}$, the copies of $F$ in $\mathcal{C}$ are in 1-1 correspondence with lifts of $\widehat{F}$ in $\widetilde{K}$, where $\widehat{F}$ is dual to $F$. Now we identify an element $F_{1}$ as a face of another element $F_{2}$ of $\mathcal{C}$, if the subcomplex of $\widetilde{K}$ associated with $F_{2}$ is contained in the subcomplex of $\widetilde{K}$ associated with $F_{1}$. Then we obtained $\mathbb{D}_{\mathcal{A}}$ from $\mathcal{C}$ after all such identifications. In particular, there is natural map $\mathbb{D}_{\mathcal{A}}\to D_{\mathcal{A}}$, coming from quotienting the action of $\mathbb{D}_{\mathcal{A}}$ by the action of $\pi_{1}\mathcal{U}$.
@@
Note that when $\mathcal{A}$ is a central arrangement, the space $\mathbb{D}_{\mathcal{A}}$ is not very interesting. In such case, there is another complex, called the spherical Deligne complex, defined as follows. Let $S_{\mathcal{A}}$ be the unit sphere, endowed with the polyhedron complex structure coming the the intersection of the unit sphere with $\mathcal{A}$. Then there is 1-1 correspondence between open cells in $S_{\mathcal{A}}$ and elements in $\operatorname{Fan}(\mathcal{A})$ which are not 0-dimensional. A face $F$ of $S_{\mathcal{A}}$ is dual to a face $F^{\prime}$ of $\Sigma_{\mathcal{A}}$ if the barycenter of $F^{\prime}$ is contained in the fan associated with $F$. In this case, we also say $F$ is dual to the standard subcomplex $\widehat{F}^{\prime}$ of $\widehat{\Sigma}_{\mathcal{A}}$. This gives a complex of group structure on $S_{\mathcal{A}}$ as before, where the local group on $F$ is the fundamental group of $\widehat{F}^{\prime}$. Then the spherical Deligne complex for the central arrangement $\mathcal{A}$, denoted $\mathbb{S}\mathbb{D}_{\mathcal{A}}$, is defined to the development complex of this complex of group over $S_{\mathcal{A}}$. This gives a natural map $\mathbb{S}\mathbb{D}_{\mathcal{A}}\to S_{\mathcal{A}}$. The complex $\mathbb{S}\mathbb{D}_{\mathcal{A}}$ also has a similar alternative description in terms of lifts of standard complexes of $\widehat{\Sigma}_{\mathcal{A}}$ in the universal cover as in the previous paragraph.
@@
Let $\mathcal{A}$ be an arbitrary affine arrangement. Let $x$ be a vertex in $D_{\mathcal{A}}$. The local arrangement of $\mathcal{A}$ at $x$, denoted by $\mathcal{A}_{x}$, is defined to be the collection of elements of $\mathcal{A}$ that contain $x$. Note that $\mathcal{A}_{x}$ is an central arrangement, and $\operatorname{lk}(x,D_{\mathcal{A}})$ can be naturally identified as a subcomplex of $S_{\mathcal{A}_{x}}$. The following is consequence of the description of $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{S}\mathbb{D}_{\mathcal{A}_{x}}$.
@@
Let $x^{\prime}\in\mathbb{D}_{\mathcal{A}}$ be a vertex which is mapped to $x\in D_{\mathcal{A}}$ under $\mathbb{D}_{\mathcal{A}}\to D_{\mathcal{A}}$. Let $N$ be the inverse image of $\operatorname{lk}(x,D_{\mathcal{A}})$ (viewed as a subset of $S_{\mathcal{A}_{x}}$) under the map $\mathbb{S}\mathbb{D}_{\mathcal{A}_{x}}\to S_{\mathcal{A}_{x}}$. Then $\operatorname{lk}(x^{\prime},\mathbb{D}_{\mathcal{A}})\cong N$.
@@ 2.5 Collapsing hyperplanes
Let $\mathcal{A}$ be an arrangement of affine hyperplanes in $V$. Let $\mathcal{A}^{\prime}\subset\mathcal{A}$ be a sub-collection of hyperplanes. Then there is a cellular map $\widehat{\Sigma}_{\mathcal{A}}\to\widehat{\Sigma}_{\mathcal{A}^{\prime}}$ defined as follows. First we define a map $\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{A}^{\prime}}$. Note that each fan of $\mathcal{A}$ is contained in a unique fan of $\mathcal{A}^{\prime}$. As vertices of the barycentric subdivision $b\Sigma_{\mathcal{A}}$ of $\Sigma_{\mathcal{A}}$ are in one to one correspondence with fans of $\mathcal{A}$, this gives a map from the vertex set of $b\Sigma_{\mathcal{A}}$ to the vertex set of $b\Sigma_{\mathcal{A}}^{\prime}$. One readily checks that vertices of a simplex are mapped to vertices of another simplex. So we can extend linearly to obtain a map $c:\Sigma_{\mathcal{A}}\cong b\Sigma_{\mathcal{A}}\to b\Sigma_{\mathcal{A}^{\prime}}\cong\Sigma_{\mathcal{A}^{\prime}}$. Note that
@@
    $c$ maps a face of $\Sigma_{\mathcal{A}}$ onto a face of $\Sigma_{\mathcal{A}^{\prime}}$;
@@
    for a face $F\subset\Sigma_{\mathcal{A}}$ and two vertices $v,v^{\prime}$ of $\Sigma_{\mathcal{A}}$ satisfying $\operatorname{Proj}_{F}(v^{\prime})=\operatorname{Proj}_{F}(v)$, we have $\operatorname{Proj}_{E}(c(v^{\prime}))=\operatorname{Proj}_{E}(c(v))$ where $E=c(F)$.
Thus $c:\Sigma_{\mathcal{A}}\cong b\Sigma_{\mathcal{A}}\to b\Sigma_{\mathcal{A}^{\prime}}\cong\Sigma_{\mathcal{A}^{\prime}}$ induces a continuous map $\hat{c}:\widehat{\Sigma}_{\mathcal{A}}\to\widehat{\Sigma}_{\mathcal{A}^{\prime}}$.
The map $\hat{c}$ restricted to $\widehat{\Sigma}_{\mathcal{A}}^{(1)}$ has a more straightforward description. As each component of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}}H$ lies in a unique component of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}^{\prime}}H$, this gives $\hat{c}:\widehat{\Sigma}_{\mathcal{A}}^{(0)}\to\widehat{\Sigma}_{\mathcal{A}^{\prime}}^{(0)}$. Moreover, two adjacent components of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}}H$ are either contained in the same component of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}^{\prime}}H$ or correspond to two adjacent components of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}^{\prime}}H$. This gives $\hat{c}:\widehat{\Sigma}_{\mathcal{A}}^{(1)}\to\widehat{\Sigma}_{\mathcal{A}^{\prime}}^{(1)}$, where an edge of $\widehat{\Sigma}_{\mathcal{A}}^{(1)}$ is collapsed to a single point if its endpoints are sent to the same point of $\widehat{\Sigma}_{\mathcal{A}^{\prime}}^{(1)}$.

### v3

@@ 4.2 Salvetti complex
The following is a construction originated in [GP12].
@@
In the rest of this subsection, we look at the case when $W_{S}$ is a finite Coxeter group with its canonical representation $\rho:W_{S}\to GL(n,\mathbb{R})$ [Dav12]. Recall a reflection in $W_{S}$ is a conjugate of a standard generator of $W_{S}$. Each reflection fixes a hyperplane in $\mathbb{R}^{n}$, which we call a reflection hyperplane. Let $\mathcal{A}$ be the collection of all reflection hyperplanes in $\mathbb{R}^{n}$. The hyperplane arrangement $\mathcal{A}$ is called the reflection arrangement associated with $W_{S}$. The following facts are standard, see e.g. [Par14].
@@
    The fundamental group $\pi_{1}(M(\mathcal{A}\otimes\mathbb{C}))$ is isomorphic to the pure Artin group $PA_{S}$ [vdL83];
@@
    As the action of $W_{S}$ permutes elements in $\mathcal{A}$, there is an induced action $W_{S}\curvearrowright M(\mathcal{A}\otimes\mathbb{C})$ and an induced action $W_{S}\curvearrowright\widehat{\Sigma}_{\mathcal{A}}$ which are free. The quotient of each of these two spaces have $\pi_{1}$ isomorphic to $A_{S}$.
@@
    The 2-skeleton of $\widehat{\Sigma}_{\mathcal{A}}/W_{S}$ is isomorphic to the presentation complex of $A_{S}$.
@@
When $\mathcal{A}$ is the reflection arrangement associated with a finite Coxeter group $W_{S}$, we will also write $\Sigma_{\mathcal{A}}$ and $\widehat{\Sigma}_{\mathcal{A}}$ as $\Sigma_{S}$ and $\widehat{\Sigma}_{S}$. Note that the 1-skeleton of $\widehat{\Sigma}_{\mathcal{A}}$ is isomorphic to the Cayley graph of $W_{S}$, and the 1-skeleton of $\Sigma_{S}$ is isomorphic to the unoriented Cayley graph of $W_{S}$ (by smashing each double edge of the usual Cayley graph to a single edge). Thus edges of $\widehat{\Sigma}_{\mathcal{A}}$ and $\Sigma_{\mathcal{A}}$ are labeled by elements of $S$. Let $K$ be a subset, or an edge path in $\widehat{\Sigma}_{\mathcal{A}}$ or $\Sigma_{\mathcal{A}}$. Then $\operatorname{Supp}(K)$ is defined to the collection of vertices of $S$ which appear as the label of an edge which is contained in $K$. The type of a standard subcomplex of $\Sigma_{S}$ or $\widehat{\Sigma}_{S}$ is defined to be the support of this subcomplex.
Note that the Coxeter complex $\mathfrak{C}_{S}$ of $W_{S}$ is homeomorphic to a sphere. More precisely, elements in $\mathcal{A}$ cuts the unit sphere of $\mathbb{R}^{n}$ into a simplicial complex, which is isomorphic to $\mathfrak{C}_{S}$. From this, we know that $\mathfrak{C}_{S}$ and $\Sigma_{\mathcal{A}}$ are dual complexes of each other. We record the following description of the Artin complex $\Delta_{S}$ in terms of $\widehat{\Sigma}_{\mathcal{A}}$, which will be used later.
Remark 4.9.
Let $X$ be the universal cover of $\widehat{\Sigma}_{\mathcal{A}}$. A lift of a standard subcomplex in $\widehat{\Sigma}_{\mathcal{A}}$ is a connected component of the inverse image of this subcomplex under the map $X\to\widehat{\Sigma}_{\mathcal{A}}$. Vertices of $\Delta_{S}$ are in 1-1 correspondence with lifts standard subcomplexes of $\widehat{\Sigma}_{\mathcal{A}}$ of type $\hat{s}$ for some $s\in S$. A collection of vertices span a simplex if their associated lifts have non-trivial common intersection.
@@ 4.3 Falk complexes for affine arrangements
Let $\mathcal{A}$ be a real affine arrangement. Let $D_{\mathcal{A}}$ be the union of all elements in $\operatorname{Fan}(\mathcal{A})$ which are bounded in $\mathbb{R}^{n}$. Note that $D_{\mathcal{A}}$ is polyhedron complex, whose open cells are elements in $\operatorname{Fan}(\mathcal{A})$ that are contained in $D_{\mathcal{A}}$. A face of $D_{\mathcal{A}}$ is defined to be the closure of an open cell of $D_{\mathcal{A}}$. Each face of $D_{\mathcal{A}}$ is a disjoint union of fans. Top dimensional faces in $D_{\mathcal{A}}$ are in 1-1 correspondence with components of $\mathbb{R}^{n}\setminus\cup_{H\in\mathcal{A}}H$ which are bounded. A face $F$ of $D_{\mathcal{A}}$ is dual to a face $F^{\prime}$ of $\Sigma_{\mathcal{A}}$ if the barycenter of $F^{\prime}$ (as defined in the beginning of Section 4.1) is contained in the interior of $F$. In this case, we will also say $F\subset D_{\mathcal{A}}$ is dual to the standard subcomplex $\widehat{F}^{\prime}$ of $\widehat{\Sigma}_{\mathcal{A}}$.
@@
We define a simple complex of group structure $\mathcal{U}$ on $D_{\mathcal{A}}$ as follows (see [BH99, Definition II.12.11]). For a face $F$ of $D_{\mathcal{A}}$, the local group at $F$ is defined to be the fundamental group of the standard subcomplex of $\widehat{\Sigma}_{\mathcal{A}}$ which is dual to $F$. The morphisms between the local groups are induced by inclusions of the associated subcomplexes. By Lemma 4.11, all the morphisms between the local groups are injective, hence $\mathcal{U}$ is a simple complex of groups. Moreover, it follows from the retraction in Definition 4.6 that each local group injects into $\pi_{1}(\widehat{\Sigma}_{\mathcal{A}})$. Thus $\mathcal{U}$ is developable with $\pi_{1}(\widehat{\Sigma}_{\mathcal{A}})=\pi_{1}\mathcal{U}$.
@@
Let $\mathbb{D}_{\mathcal{A}}$ be the development complex of $\mathcal{U}$ (see [BH99, Theorem II.12.18]). Then $\mathbb{D}_{\mathcal{A}}$ is simply-connected ([BH99, Corollary II.12.21]). The complex $\mathbb{D}_{\mathcal{A}}$ is defined to be the Falk complex of the affine arrangement $\mathcal{A}$, motivated by the work of Falk [Fal95]. We now give an alternative description of $\mathbb{D}_{\mathcal{A}}$. Let $\widetilde{K}$ be the universal cover of $\widehat{\Sigma}_{\mathcal{A}}$. We start with a disjoint union of a collection $\mathcal{C}$ of multiple copies of faces of $D_{\mathcal{A}}$ as follows: for each face $F$ of $D_{\mathcal{A}}$, the copies of $F$ in $\mathcal{C}$ are in 1-1 correspondence with lifts of $\widehat{E}$ in $\widetilde{K}$, where $\widehat{E}$ is dual to $F$. Now we identify an element $F_{1}$ as a face of another element $F_{2}$ of $\mathcal{C}$, if the subcomplex of $\widetilde{K}$ associated with $F_{2}$ is contained in the subcomplex of $\widetilde{K}$ associated with $F_{1}$. Then we obtained $\mathbb{D}_{\mathcal{A}}$ from $\mathcal{C}$ after all such identifications. In particular, there is natural map $\mathbb{D}_{\mathcal{A}}\to D_{\mathcal{A}}$, coming from quotienting $\mathbb{D}_{\mathcal{A}}$ by the action of $\pi_{1}\mathcal{U}$.
@@
When $\mathcal{A}$ is a central arrangement, there is another associated complex, called the spherical Deligne complex, defined as follows. Let $S_{\mathcal{A}}$ be the unit sphere, endowed with the polyhedron complex structure coming the the intersection of the unit sphere with $\mathcal{A}$. Then there is 1-1 correspondence between open cells in $S_{\mathcal{A}}$ and elements in $\operatorname{Fan}(\mathcal{A})$ which are not 0-dimensional. A face $F$ of $S_{\mathcal{A}}$ is dual to a face $F^{\prime}$ of $\Sigma_{\mathcal{A}}$ if the barycenter of $F^{\prime}$ is contained in the fan associated with $F$. In this case, we also say $F$ is dual to the standard subcomplex $\widehat{F}^{\prime}$ of $\widehat{\Sigma}_{\mathcal{A}}$. This gives a complex of group structure on $S_{\mathcal{A}}$ as before, where the local group on $F$ is the fundamental group of $\widehat{F}^{\prime}$. Then the spherical Deligne complex for the central arrangement $\mathcal{A}$, denoted $\mathbb{S}_{\mathcal{A}}$, is defined to the development complex of this complex of group over $S_{\mathcal{A}}$. This gives a natural map $\mathbb{S}_{\mathcal{A}}\to S_{\mathcal{A}}$. The complex $\mathbb{S}_{\mathcal{A}}$ also has a similar alternative description in terms of lifts of standard complexes of $\widehat{\Sigma}_{\mathcal{A}}$ in the universal cover as in the previous paragraph. The complex $\mathbb{S}_{\mathcal{A}}$ is simply-connected if $S_{\mathcal{A}}$ is simply-connected, i.e. $n\geq 3$.
@@
Let $\mathcal{A}$ be an arbitrary affine arrangement. Let $x$ be a vertex in $D_{\mathcal{A}}$. The local arrangement of $\mathcal{A}$ at $x$, denoted by $\mathcal{A}_{x}$, is defined to be the collection of elements of $\mathcal{A}$ that contain $x$. Note that $\mathcal{A}_{x}$ is an central arrangement, and $\operatorname{lk}(x,D_{\mathcal{A}})$ can be naturally identified as a subcomplex of $S_{\mathcal{A}_{x}}$. The following is consequence of the description of $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{S}_{\mathcal{A}_{x}}$.
@@
Let $x^{\prime}\in\mathbb{D}_{\mathcal{A}}$ be a vertex which is mapped to $x\in D_{\mathcal{A}}$ under $\mathbb{D}_{\mathcal{A}}\to D_{\mathcal{A}}$. Let $N$ be the inverse image of $\operatorname{lk}(x,D_{\mathcal{A}})$ (viewed as a subset of $S_{\mathcal{A}_{x}}$) under the map $\mathbb{S}_{\mathcal{A}_{x}}\to S_{\mathcal{A}_{x}}$. Then $\operatorname{lk}(x^{\prime},\mathbb{D}_{\mathcal{A}})\cong N$.
@@ 4.4 Collapsing hyperplanes
Let $\mathcal{A}$ be an arrangement of affine hyperplanes in $\mathbb{R}^{n}$. Let $\mathcal{A}^{\prime}\subset\mathcal{A}$ be a sub-collection of hyperplanes. Then there is a cellular map $\widehat{\Sigma}_{\mathcal{A}}\to\widehat{\Sigma}_{\mathcal{A}^{\prime}}$ defined as follows. First we define a map $\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{A}^{\prime}}$. Note that each fan of $\mathcal{A}$ is contained in a unique fan of $\mathcal{A}^{\prime}$. As vertices of the barycentric subdivision $b\Sigma_{\mathcal{A}}$ of $\Sigma_{\mathcal{A}}$ are in one to one correspondence with fans of $\mathcal{A}$, this gives a map from the vertex set of $b\Sigma_{\mathcal{A}}$ to the vertex set of $b\Sigma_{\mathcal{A}}^{\prime}$. One readily checks that vertices of a simplex are mapped to vertices of another simplex. So we can extend linearly to obtain a map $\kappa:b\Sigma_{\mathcal{A}}\to b\Sigma_{\mathcal{A}^{\prime}}$, which can also be viewed as a map $\kappa:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{A}^{\prime}}$. Note that
@@
    $\kappa$ maps a face of $\Sigma_{\mathcal{A}}$ onto a face of $\Sigma_{\mathcal{A}^{\prime}}$;
@@
    for a face $F\subset\Sigma_{\mathcal{A}}$ and two vertices $v,v^{\prime}$ of $\Sigma_{\mathcal{A}}$ satisfying $\operatorname{Proj}_{F}(v^{\prime})=\operatorname{Proj}_{F}(v)$, we have $\operatorname{Proj}_{E}(\kappa(v^{\prime}))=\operatorname{Proj}_{E}(\kappa(v))$ where $E=\kappa(F)$.
Thus $\kappa:\Sigma_{\mathcal{A}}\cong b\Sigma_{\mathcal{A}}\to b\Sigma_{\mathcal{A}^{\prime}}\cong\Sigma_{\mathcal{A}^{\prime}}$ induces a continuous map $\hat{\kappa}:\widehat{\Sigma}_{\mathcal{A}}\to\widehat{\Sigma}_{\mathcal{A}^{\prime}}$.
The map $\hat{\kappa}$ restricted to $\widehat{\Sigma}_{\mathcal{A}}^{(1)}$ has a more straightforward description. As each component of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}}H$ lies in a unique component of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}^{\prime}}H$, this gives $\hat{\kappa}:\widehat{\Sigma}_{\mathcal{A}}^{(0)}\to\widehat{\Sigma}_{\mathcal{A}^{\prime}}^{(0)}$. Moreover, two adjacent components of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}}H$ are either contained in the same component of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}^{\prime}}H$ or correspond to two adjacent components of $\mathbb{R}^{n}-\cup_{H\in\mathcal{A}^{\prime}}H$. This gives $\hat{\kappa}:\widehat{\Sigma}_{\mathcal{A}}^{(1)}\to\widehat{\Sigma}_{\mathcal{A}^{\prime}}^{(1)}$, where an edge of $\widehat{\Sigma}_{\mathcal{A}}^{(1)}$ is collapsed to a single point if its endpoints are sent to the same point of $\widehat{\Sigma}_{\mathcal{A}^{\prime}}^{(1)}$.